OpenStudy (anonymous):

simplify: x(1-2x)^(-3/2)+(1-2x)^(-1/2) PLEASE SHOW STEPS!!

OpenStudy (anonymous):

this looks like a derivative?

OpenStudy (anonymous):

if you really want to simplify, it is probably best to get rid of the negative exponents so we can see what this is

OpenStudy (anonymous):

\[\frac{x}{\sqrt{(1-2x)^3}}+\frac{1}{\sqrt{1-2x}}\] is what you really have and if you want to add them you have to give them a common denominator that would be \[\sqrt{1-2x}^3\] and you do that by multiplying the second fraction top and bottom by \[1-2x\]

OpenStudy (anonymous):

so you get \[\frac{x}{\sqrt{(1-2x)^3}}+\frac{1}{\sqrt{1-2x}}\times \frac{1-2x}{1-2x}\] \[=\frac{x+1-2x}{\sqrt{(1-2x)^3}}=\frac{1-x}{\sqrt{(1-2x)^3}}\]\]